Using prices in 2020 as a base year, calculate real GDP in 2020 and 2021.

Q 2020         price 2020 Q 2021  price 2021

Apples                   1          240               3

oranges                 2           90               4

plams                    4          120               5 Output Quantity (Q) and prices in a simply economy


1
Expert's answer
2021-01-31T08:29:02-0500

Let's calculate the real GDP in 2020 year:


"Real\ GDP_{2020}=120\cdot\$1+65\cdot\$2+80\cdot\$4=\$570."

Let's calculate the real GDP in 2021 year:


"Real\ GDP_{2021}=240\cdot\$1+90\cdot\$2+120\cdot\$4=\$900."